ROTATIONAL SPECTRUM OF THE CO_{2}–CO_{2}–OCS TRIMER

\underline{SEAN A. PEEBLES}, ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, Ann Arbor, MI 48109-1055, USA.

The rotational spectrum of the CO_{2}–CO_{2}–OCS trimer has been observed for the first time by pulsed supersonic nozzle Fourier-transform microwave spectroscopy. 52 observed transitions for the normal species were fit to a Watson Areduction Hamiltonian to obtain the following spectroscopic constants: A = 1299.1386(4)MHz, B = 973.7411(3)MHz, C = 648.6719(3)MHz, Δ_{J} = 1.270(5)kHz, Δ_{JK} = 2.41(2)kHz, δ_{J} = 0.336(2)kHz, δ_{K} = 1.95(1)kHz.

Measurements of the rotational spectra of a further eight isotopic species are consistent with a non-planar, barrel-like structure for the trimer which aligns the three monomer axes roughly parallel to one another. An analysis of the spectroscopic data will be presented and compared with results obtained from theoretical modeling of this system using atom-atom parameters.